A divalent cation is isoelctronic of CO2 and has (Z+2) neutrons . What number of molecules of its carbonate are present in 55 grams of metal carbonate ?
 

Grade:9

1 Answers

Piyush Kumar Behera
7 years ago
As the it is a divalent cation ,let it be X2+.
Number of electrons in COis 6+8+8=22 electrons.
As X2+ is isoelectronic with CO2 .So number of electrons in it is 22.
Number of electrons in its neutral atom is 22+2=24(Because X had lost electron to become X2+).Hence the atomic number of X is 24.
number of protons(Z)=atomic number =24
number of neutrons=Z+2=26
Mass Number=Z+Z+2=50u
Formula of its carbonate= XCO3
Molecurar mass of its carbonate=50+12+16 x 3=110
Hence 1mole=110 g=6.023x1023
=>  55g=6.023 x 1023/2 =3.011x1023
Hence the number of molecule is 3.011x1023